An epidemiologist wished to compare two rabies vaccines. Subjects who had previously received rabies vaccine were...

  1. An epidemiologist wished to compare two rabies vaccines. Subjects who had previously received rabies vaccine were divided in two groups. Group 1 received a booster does of type 1 vaccine, and Group 2 received a booster dose of type 2 vaccine. Antibody responses were recorded two weeks later. The means, standard deviations, and sample sizes for the two groups were as follows:

Group

n

(mean) x

s

    1

10

4.5

2.5

    2

9

2.5

2.0

The data come from normally distributed populations. Do these data indicate that there is a difference in the effectiveness of the two vaccine used for booster doses? Let the level of significance = 0.01
Answer #1


Answer Date: 05/11/2019 To test the hypothesis is that there is a difference in the effectiveness of the two vaccines used foThe t test-statistics is, +- (T1 - 12)-(4- ) (4.50 – 2.50)-0 5.191 5.191 V 10 9 = 1.910 The t test-statistics is 1.910. The td.f.=n +n2 - 2 = 10 + 9-2 =17 The t-critical value is - 2.898 and 2.898). Decision The statistical decision is that the t-tes
Answer #2
23, 26 Dhsu Ei3ri 3fnek
source: Biostatistics
answered by: Odmayie